OpenStudy (rollfaceblah):

The length of the hypotenuse of a right triangle is 24 inches. If the length of one leg is 8 inches, what is the approximate length of the other leg? 16.0 inches 18.4 inches 22.6 inches 25.3 inches

OpenStudy (welshfella):

Use pythagoras theorem

OpenStudy (michele_laino):

hint: the requested side is given by the subsequent computation: \[\Large \sqrt {{{24}^2} - {8^2}} = \sqrt {576 - 64} = ...?\]

OpenStudy (rollfaceblah):

i got 22.6274... ?